OpenStudy (anonymous):

Really need help with 3 questions! The correct simplification of the expression -6x^2(4x - 7x^2 - 2)? The correct simplification of the expression (-2x + 6y)(4x - 6y)? The correct simplification of the expression (3x - 6)(2x^2 - 4x - 5)?

OpenStudy (hoblos):

a(b+c) = ab + ac (a+b)(c+d) = ac + ad + bc + bd

OpenStudy (anonymous):

also you'll need.. \[\LARGE x^a\cdot x^b=x^{a+b}\]

OpenStudy (anonymous):

Can you tell me the answer please? This is for a study guide and if I fail the test, I have to do Algebra again

OpenStudy (anonymous):

well you have to try at least, show us what you've done so far, then we'll correct your mistakes. Deal ? ;)

OpenStudy (anonymous):

Okay hold on

OpenStudy (anonymous):

I know the first one deals with the Power of a Power rule so it should equal 42x^4 + 24x^3 - 12x^2. The second one is the Products of Powers so it should be -8x^2 + 36xy - 36y^2. The last one is the one I am having a lot of trouble with. I am terrible at this :(

OpenStudy (anonymous):

the first one is close just look again signs minus and plus, ... the second one it's correct well done.

OpenStudy (anonymous):

Okay here is the oter answer I got for the first -42x^4 + 24x^3 - 12x^2

OpenStudy (anonymous):

\[\Large -6x^2(4x-7x^2-2)=-24x^3+42x^4+12x^2\] because ... \[\LARGE (-)\cdot (-)=+\]

OpenStudy (anonymous):

that was the first one... THE SECOND ONE YOU HAVE IT CORRECT ... now I'll give you a hint on the third one .

OpenStudy (anonymous):

I know my last answer is wrong 6x^3 + 9x + 30

OpenStudy (anonymous):

is this your answer for the third question ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hmmm ... hold on

OpenStudy (anonymous):

\[\Large (3x-6)(2x^2-4x-5)=\] \[\Large =6x^3-12x^2-15x-12x^2+24x+30\] can you continue ?

OpenStudy (anonymous):

ohhh! I see what I did wrong! Thanks!

OpenStudy (anonymous):

... Give me your final answer , I want to be sure it's correct ! ;)

OpenStudy (anonymous):

okay give me a sec

OpenStudy (anonymous):

6x^3 - 24x^2 + 39x + 30?

OpenStudy (anonymous):

why 39x ?? ... look again signs carefully !

OpenStudy (anonymous):

can you find the mistake ?

OpenStudy (anonymous):

Aww can't! I have been working on this study guide all night

OpenStudy (anonymous):

-15x+24x= ?

OpenStudy (anonymous):

ohhh i see 6x^3 - 24x^2 + 9x + 30

OpenStudy (anonymous):

That's correct. WELL DONE ! ;) ...

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

Glad to help, and you're welcome ;)